OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x – 6y = –15 x – 3y = –3

OpenStudy (texaschic101):

x = 6y - 15 sub 6y - 15 in for x in second equation x - 3y = -3 (6y - 15) - 3y = -3 3y - 15 = -3